### v1

@@ Abstract
(a) Ground truth, 0-1 loss: 0, KK=5, dd=3
(b) CART, 0-1 loss: 56, KK=15, dd=5
(c) ConTree, 0-1 loss: 34, KK=15, dd=4
(d) HODT (Our), 0-1 loss: 18, KK=3, dd=3
Figure 1: Synthetic dataset and decision trees with corresponding 0-1 loss, tree sizes ($K$), and tree depths ($d$).
@@ 1 Introduction
By contrast, decision trees that allow hyperplane splits are often both smaller and more accurate. When the underlying decision regions involve complex or non-convex boundaries, trees with hyperplane or hypersurface splitting rules provide a more faithful representation, whereas axis-parallel trees remain limited to hyper-rectangular partitions. For instance, Figure LABEL:fig:_comparision shows a synthetic dataset generated from three quadratic hypersurfaces (conic sections). The decision regions produced by classical CART and by the state-of-the-art optimal decision tree algorithm ConTree (Brița et al., 2025), displayed in the two middle panels, are constrained to irregular axis-aligned rectangles. Both CART and ConTree yield trees with 15 splitting rules, of depths 5 and 4, respectively.
The above example motivates the design of algorithms for constructing decision trees with more flexible splitting rules, such as hyperplanes or, more generally, hypersurfaces. However, to the best of our knowledge, no algorithm—whether exact, approximate, or heuristic—has been proposed for learning decision trees with hypersurface splitting rules. In this second paper of the series, we present the first such algorithm, building on the results from Part I to solve the size-constrained optimal hypersurface decision tree (HODT) problem. The algorithm is developed through several key results, including the prefix-closed properties of crossed hyperplanes and the incremental generation of ancestry relation matrices. The main contributions of this paper are as follows:
@@
The structure of this series is as follows. Part I introduced the algorithmic and geometric foundations of the ODT problem. Part II, presented here, develops the first algorithm for solving the HODT problem, building directly on the foundations of Part I. Section 3, details the construction of the algorithm: Subsection 3.1 introduces ancestry relations for hyperplanes, identifying a critical property—crossed hyperplanes—and proving that any set of hyperplanes containing such a pair cannot form a proper decision tree. This result forms the basis for an efficient ancestry-matrix generator free of crossed hyperplanes, described in Subsection 3.2. The complete HODT algorithm is then presented in Subsection 3.3.
@@ 2 Review of Part I
In Part I, we formalized the decision tree problem through a set of axioms, referred to as the proper decision tree axiom. Based on these axioms, we derived four generic definitions of the decision tree problem, expressed unambiguously as four recursive programs. Through straightforward derivations, we proved both the existence and non-existence of dynamic programming (DP) solutions for these definitions under certain objective functions. We then examined the feasibility of solving the ODT problem under depth or size constraints, and argued that size-constrained trees are more suitable for problems with formidable combinatorial complexity. This suitability arises from their inherent parallelizability, achieved by factoring size-constrained trees via $K$-combinations.
@@
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right]$
@@
In recalling the general framework for solving the size-constrained ODT problem, an astute reader may notice that the program $\mathit{odt}_{\text{size}}$ operates on a list of splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$ rather than directly on the data. This design choice reflects our aim of providing a generic, simple, and modular framework, since decision trees are defined in terms of splitting rules rather than data instances. Consequently, when applying the framework to practical ODT problems in machine learning, a separate procedure $\mathit{gen}_{\mathit{splits}}$ must be introduced to generate the splitting rules, with its definition depending on the specific task. This modularity also provides a further advantage: adapting the framework to different problem settings requires only minimal code modifications.
@@
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M}$
where $\mathit{embed}_{M}:\left[\mathbb{R}^{D}\right]\to\left[\mathbb{R}^{G}\right]$ is the Veronese embedding ($\rho_{M}$) introduced in Part I, mapping data points $\mathit{xs}$ in $\mathbb{R}^{D}$ to the embedded dataset $\rho_{M}\left(\mathit{xs}\right)$ in $\mathbb{R}^{G}$, with $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$. Moreover, $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ can be unified into a single program $\mathit{nestedCombs}\left(K,G\right)$, for which the implementation was given in Part I. For completeness, we include its pseudocode in Algorithm 5 in Appendix A.
@@ 3.1 Ancestry relation and crossed hyperplanes
Axiom 4 of proper decision trees is highly problem-dependent. For axis-parallel decision trees, the splitting rule $h^{0}$ can be characterized by a single data item. Assuming no two data points lie on the same axis-parallel hyperplane (i.e., the set of candidate splitting rules contains no duplicates—duplicates, if present, can simply be removed without loss of generality), it follows that the relation $h_{i}^{0}\overline{\left(\swarrow\vee\searrow\right)}h_{j}^{0}$ is impossible. This is because any data point always lies on exactly one side of an axis-parallel hyperplane. Thus, for any pair of axis-parallel hyperplanes $h_{i}^{0}$ and $h_{j}^{0}$ either $h_{i}^{0}\swarrow h_{j}^{0}$ or $h_{i}^{0}\searrow h_{j}^{0}$ holds. In other words, any axis-parallel hyperplane $h_{j}^{0}$ must lie in either the left or right subregion determined by another axis-parallel hyperplane $h_{i}^{0}$.
@@
Proof.
We prove this by contradiction. Assume there exists a third hypersurface $h_{k}$ such that $h_{i}$ and $h_{j}$ lie on the left and right branches of $h_{k}$, respectively. Then $h_{k}$ must also separate the data points defining $h_{i}$ and $h_{j}$ in two decision regions. By the definition of convex combinations in Euclidean space, for any set of points $xs=\left\{\boldsymbol{x}_{1},\boldsymbol{x}_{2}\ldots\boldsymbol{x}_{D}\right\}$ on the same side of a hyperplane, i.e., $\boldsymbol{w}^{T}\boldsymbol{x}_{i}\geq 0$ or $\boldsymbol{w}^{T}\boldsymbol{x}_{i}<0$, $\forall x_{i}\in xs$, the convex combination $\boldsymbol{y}=\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}$ will satisfy $\boldsymbol{w}^{T}\boldsymbol{y}=\boldsymbol{w}^{T}\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}\geq 0$, where $\lambda_{i}\geq 0$. Thus, the convex hull of points on one side of a hyperplane remains entirely on that side.
Therefore, if $h_{k}$ separate the defining data points of $h_{i}$ and $h_{j}$ it must also separate their respective convex hulls. However, by the definition of crossed hypersurfaces, the defining points of $h_{i}$ lies on both side of $h_{j}$, and the defining points of $h_{j}$ lies on both sides of $h_{i}$. Consequently, the convex hulls of $h_{i}$ and $h_{j}$ intersect, making it impossible for a single hyperplane $h_{k}$ to separate them. This contradicts the assumption that such $h_{k}$ exists. Hence, no hypersurface can separate crossed hypersurfaces, and any combination containing them cannot form a proper decision tree. ∎
Figure 4: An example illustrating the proof of Fact 1. Demonstrating that the data items $a$, $b$, which define $h_{i}$, and $c$, $d$, which define $h_{j}$ cannot be classified into the disjoint regions defined by a third hyperplane (red).
@@
Proof.
By definition, if a pair $h_{i}$ and $h_{j}$ in $\mathit{hs}_{K}$ are crossed, then neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ nor $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable, and therefore $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$. Conversely, if $\mathit{hs}_{K}$ contains no crossed hyperplanes, then for any pair $h_{i},h_{j}\in\mathit{hs}_{K}$, either a mutual ancestry or an asymmetrical ancestry exists. Hence, a proper decision tree can be constructed. ∎
@@ 3.2 A generic, incremental, crossed-hyperplane free, ancestry relation matrix generator
[…]
@@
Proof.
We prove the fusion theorem by following reasoning
@@
The equivalence between filter fusion condition (10) and prefix-closed property (11) is straight forwards. A configuration $a:\mathit{as}$ survived in $\mathit{filter}_{p}$ will also survived in $\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(a,\mathit{filter}_{p}\left(\left[\mathit{as}\right]\right)\right)\right)$ because $p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$. ∎
@@
  1. 1.
    Input: $r_{j}$: the index of the new splitting rule; $\mathit{ncsr}$: a $M\times\left(k+1\right)\times k$ tensor, which contains a vector of $\mathit{ncr}:NCR$ configurations; $\mathit{css}$: a matrix for storing all $G$-combinations of data points of all splitting rules; $\mathit{asgn}^{\pm}$: the positive/negative predictions of each splitting rule
  2. 2.
    Output: The updated nested combinations $\mathit{ncsr}^{\prime}:\mathit{NCRs}$ without crossed hyperplanes, which is a a vector of $\left(k+2\right)\times\left(k+1\right)$ matrices
  3. 3.
    $\mathit{ncrs}^{\prime}=\left[\>\right]$
  4. 4.
    for $\mathit{ncr}\in\mathit{ncrs}$ do:
  5. 5.
    for $r_{i}\in\mathit{ncr}\left[0\right]$ do:
  6. 6.
       $p_{1}=\text{True }\text{if }\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{+}\left[r_{i}\right]\vee\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{-}\left[r_{i}\right]$ else False // the $G$-combination of data points $\mathit{css}\left[G\right]\left[r_{j}\right]$ all lies in the positive side or negative side of the hyperplane $r_{i}$,
  7. 7.
      $p_{2}=\text{True }\text{if }$ $\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{+}\left[r_{j}\right]\vee\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{-}\left[r_{j}\right]$ else False // the $G$-combination corresponds for defining rule $r_{i}$ all lies in the positive side or negative side of the splitting rule $r_{j}$, $\mathit{unrank}\left(r_{i}\right)$ takes an index and return all data points for defining splitting rule $r_{i}$,
  8. 8.
      if $p_{1}\vee p_{2}$ do:
  9. 9.
       $\mathit{ncr}^{\prime}=\mathit{update}_{\text{arMat}}\left(\mathit{ncr}\right)$ // update ancestry relation matrix
  10. 10.
       $\mathit{ncrs}^{\prime}+\!\!+\left[\mathit{ncr}^{\prime}\right]$
  11. 11.
      else:
  12. 12.
        break // break the inner loop because $\mathit{ncr}$ is infeasible
  13. 13.
    return $\mathit{ncrs}^{\prime}$
Algorithm 1 $\mathit{updates}_{\text{arMat}}$
@@
With these components, we can construct the incremental, crossed-hyperplane-free ancestry relation matrix generator $\mathit{\mathit{nestedCombsFA}}$. This is achieved by introducing a single additional line of code after line 12 of Algorithm using Algorithm 1 to update all feasible nested combinations while avoiding non-feasible combinations containing crossed hyperplanes. The resulting modification is minimal—just one line—and the complete pseudocode is provided in Appendix A.2, Algorithm 6.

### v2

@@ Abstract
Keywords: decision tree, dynamic programming, global optimal algorithm
Figure 1: Synthetic dataset (left) generated by degree-2 polynomials; axis-parallel decision trees learned by CART and the optimal algorithm (middle two); and hypersurface decision trees learned using our proposed algorithm (right), with corresponding misclassification counts, tree sizes, and tree depths.
@@ 1 Introduction
By contrast, decision trees that allow hyperplane splits are often both smaller and more accurate. When the underlying decision regions involve complex or non-convex boundaries, trees with hyperplane or hypersurface splitting rules provide a more faithful representation, whereas axis-parallel trees remain limited to hyper-rectangular partitions. For instance, Figure 1 shows a synthetic dataset generated from three quadratic hypersurfaces (conic sections). The decision regions produced by classical CART and by the state-of-the-art optimal decision tree algorithm ConTree (Brița et al., 2025), displayed in the two middle panels, are constrained to irregular axis-aligned rectangles. Both CART and ConTree yield trees with 15 splitting rules, of depths 5 and 4, respectively, whereas our method produce a tree with 3 splits of depth $3$.
The above example motivates the design of algorithms for constructing decision trees with more flexible splitting rules, such as hyperplanes or, more generally, hypersurfaces. However, to the best of our knowledge, no algorithm—whether exact, approximate, or heuristic—has been proposed for learning decision trees with hypersurface splitting rules. In this second paper of the series, we present the first such algorithm, building on the results from Part I (He, 2025) to solve the size-constrained optimal hypersurface decision tree (HODT) problem. The algorithm is developed through several key results, including the prefix-closed properties of crossed hyperplanes and the incremental generation of ancestry relation matrices. The main contributions of this paper are as follows:
@@
The structure of this series is as follows. Part I introduced the algorithmic and geometric foundations of the ODT problem. Part II, presented here, develops the first algorithm for solving the HODT problem, building directly on the foundations of Part I. Section 3 details the construction of the algorithm: Subsection 3.1 introduces ancestry relations for hyperplanes, identifying a critical property—crossed hyperplanes—and proving that any set of hyperplanes containing such a pair cannot form a proper decision tree. This result forms the basis for an efficient ancestry-matrix generator free of crossed hyperplanes, described in Subsection 3.2. The complete HODT algorithm is then presented in Subsection 3.3.
@@ 2 Review of Part I
In Part I (He, 2025), we formalized the decision tree problem through a set of axioms, referred to as the proper decision tree axiom. Based on these axioms, we derived four generic definitions of the decision tree problem, expressed unambiguously as four recursive programs. Through straightforward derivations, we proved both the existence and non-existence of dynamic programming (DP) solutions for these definitions under certain objective functions. We then examined the feasibility of solving the ODT problem under depth or size constraints, and argued that size-constrained trees are more suitable for problems with formidable combinatorial complexity. This suitability arises from their inherent parallelizability, achieved by factoring size-constrained trees via $K$-combinations.
@@
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right].$
@@
In recalling the general framework for solving the size-constrained ODT problem, an astute reader may notice that the program $\mathit{odt}_{\text{size}}$ operates on a list of splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$ rather than directly on the data. This design choice reflects our aim of providing a generic and modular framework, since decision trees are defined in terms of splitting rules rather than data instances. Consequently, when applying the framework to practical ODT problems in machine learning, a separate procedure $\mathit{gen}_{\mathit{splits}}$ must be introduced to generate the splitting rules, with its definition depending on the specific task. This modularity also provides a further advantage: adapting the framework to different problem settings requires only minimal code modifications.
@@
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M},$
where $\mathit{embed}_{M}:\left[\mathbb{R}^{D}\right]\to\left[\mathbb{R}^{G}\right]$ is the Veronese embedding ($\rho_{M}$) introduced in Part I, mapping data points $\mathit{xs}$ in $\mathbb{R}^{D}$ to the embedded dataset $\rho_{M}\left(\mathit{xs}\right)$ in $\mathbb{R}^{G}$, with $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$. The transformation from the second to the final equation unifies $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ into a single program $\mathit{nestedCombs}\left(K,G\right)$ as derived in Part I. For completeness, the pseudo-code for $\mathit{nestedCombs}\left(K,G\right)$ is provided in Algorithm 5.
@@ 3.1 Ancestry relation and crossed hyperplanes
Axiom 4 of proper decision trees is highly problem-dependent. For axis-parallel decision trees, the splitting rule $h^{0}$ can be characterized by a single data item. Assuming no two data points lie on the same axis-parallel hyperplane (i.e., the set of candidate splitting rules contains no duplicates—duplicates, if present, can simply be removed without loss of generality), it follows that the relation $h_{i}^{0}\overline{\left(\swarrow\vee\searrow\right)}h_{j}^{0}$ is impossible. This is because any data point always lies on exactly one side of an axis-parallel hyperplane. Thus, for any pair of axis-parallel hyperplanes $h_{i}^{0}$ and $h_{j}^{0}$ either $h_{i}^{0}\swarrow h_{j}^{0}$ or $h_{i}^{0}\searrow h_{j}^{0}$ holds. In other words, any axis-parallel hyperplane $h_{j}^{0}$ must lie in either the left or right sub-region determined by another axis-parallel hyperplane $h_{i}^{0}$.
@@
Proof We prove this by contradiction. Assume there exists a third hypersurface $h_{k}$ such that $h_{i}$ and $h_{j}$ lie on the left and right branches of $h_{k}$, respectively. Then $h_{k}$ must also separate the data points defining $h_{i}$ and $h_{j}$ in two decision regions. By the definition of convex combinations in Euclidean space, for any set of points $xs=\left\{\boldsymbol{x}_{1},\boldsymbol{x}_{2}\ldots\boldsymbol{x}_{D}\right\}$ on the same side of a hyperplane, i.e., $\boldsymbol{w}^{T}\boldsymbol{x}_{i}\geq 0$ or $\boldsymbol{w}^{T}\boldsymbol{x}_{i}<0$, $\forall x_{i}\in xs$, the convex combination $\boldsymbol{y}=\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}$ will satisfy $\boldsymbol{w}^{T}\boldsymbol{y}=\boldsymbol{w}^{T}\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}\geq 0$, where $\lambda_{i}\geq 0$. Thus, the convex hull of points on one side of a hyperplane remains entirely on that side.
Therefore, if $h_{k}$ separate the defining data points of $h_{i}$ and $h_{j}$ it must also separate their respective convex hulls. However, by the definition of crossed hypersurfaces, the defining points of $h_{i}$ lies on both side of $h_{j}$, and the defining points of $h_{j}$ lies on both sides of $h_{i}$. Consequently, the convex hulls of $h_{i}$ and $h_{j}$ intersect, making it impossible for a single hyperplane $h_{k}$ to separate them. This contradicts the assumption that such $h_{k}$ exists. Hence, no hypersurface can separate crossed hypersurfaces, and any combination containing them cannot form a proper decision tree.
Figure 4: An example illustrating the proof of Fact 2. Demonstrating that the data items $a$, $b$, which define $h_{i}$, and $c$, $d$, which define $h_{j}$ cannot be classified into the disjoint regions defined by a third hyperplane (red).
@@
Proof By definition, if a pair $h_{i}$ and $h_{j}$ in $\mathit{hs}_{K}$ are crossed, then neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ nor $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable, and therefore $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$. Conversely, if $\mathit{hs}_{K}$ contains no crossed hyperplanes, then for any pair $h_{i},h_{j}\in\mathit{hs}_{K}$, either a mutual ancestry or an asymmetrical ancestry exists. Hence, a proper decision tree can be constructed.
@@ 3.2 A generic, incremental, crossed-hyperplane free, ancestry relation matrix generator
$\mathit{calARM}$ $\left(\left[r\right]\right)=\left[\begin{array}[]{c}r\\ 0\end{array}\right]$
@@
Proof We prove the fusion theorem by following reasoning
@@
The equivalence between filter fusion condition (10) and prefix-closed property (11) is straight forwards. A configuration $a:\mathit{as}$ survived in $\mathit{filter}_{p}$ will also survived in $\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(a,\mathit{filter}_{p}\left(\left[\mathit{as}\right]\right)\right)\right)$ because $p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$.
@@
Algorithm 1 Update Ancestry Relation ($\mathit{updates}_{arMat}$)
1:Input: $\mathit{r}_{j}$: index of new splitting rule; $\mathit{ncrs}$: $M\times(k+1)\times k$ tensor of $\mathit{ncr}$ configurations; $\mathit{css}$: matrix of $G$-combinations; $\mathit{asgn}^{\pm}$: positive/negative predictions
2:Output: Updated nested combinations $\mathit{ncrs}^{\prime}$: $\mathit{NCRs}$ without crossed hyperplanes, as $(k+2)\times(k+1)$ matrices
[…]
4:for $\mathit{ncr}\in\mathit{ncrs}$ do
5:  for $\mathit{r}_{i}\in\mathit{ncr}[0]$ do
6:   $p_{1}\leftarrow\top$ if $\mathit{css}[G][\mathit{r}_{j}]\in\mathit{asgn}^{+}[\mathit{r}_{i}]\vee\mathit{css}[G][\mathit{r}_{j}]\in\mathit{asgn}^{-}[\mathit{r}_{i}]$ else $\bot$ $\triangleright$ Check if $G$-combination lies on one side of $\mathit{r}_{i}$
7:   $p_{2}\leftarrow\top$ if $\mathit{unrank}(\mathit{r}_{i})\in\mathit{asgn}^{+}[\mathit{r}_{j}]\vee\mathit{unrank}(\mathit{r}_{i})\in\mathit{asgn}^{-}[\mathit{r}_{j}]$ else $\bot$ $\triangleright$ Check if rule $\mathit{r}_{i}$ lies on one side of $\mathit{r}_{j}$
8:   if $p_{1}\vee p_{2}$ then
9:    $\mathit{ncr}^{\prime}\leftarrow\mathit{update}_{arMat}(\mathit{ncr})$ $\triangleright$ Update ancestry relation matrix
10:    $\mathit{ncrs}^{\prime}\leftarrow\mathit{ncrs}^{\prime}+[\mathit{ncr}^{\prime}]$ $\triangleright$ Append updated configuration
11:   else
12:    break $\triangleright$ Skip infeasible $\mathit{ncr}$
13:   end if
14:  end for
15:end for
16:return $\mathit{ncrs}^{\prime}$
@@
With these components, we can construct the incremental, crossed-hyperplane-free ancestry relation matrix generator $\mathit{\mathit{nestedCombsFA}}$. This is achieved by introducing a single additional line of code after line 12 of Algorithm using Algorithm 1 to update all feasible nested combinations while avoiding non-feasible combinations containing crossed hyperplanes. The resulting modification is minimal—just one line—and the complete pseudo-code is provided in Appendix A.2, Algorithm 6.